5sin^2 90deg + 2cos^2 270deg - tan 360deg
Answer by ikleyn(52946)

.
5sin^2(90°) + 2cos^2(270°) - tan(360°)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

sin(90°) = 1; cos(270°) = 0; tan(360°) = 0.

Therefore, 5sin^2(90°) + 2cos^2(270°) - tan(360°) = 5*1 + 2*0 -0 = 5.